OpenStudy (anonymous):

determine if h(x)=5/x + 2 is an inverse of f(x) = 5/(x-2)

OpenStudy (anonymous):

You can make this determination by re-writing f(x) as: f(y) = x = 5/(y - 2) and solve for "y" and see if you get the right side of h(x)

OpenStudy (anonymous):

i got 5x+2 so im assuming the answer is no?

OpenStudy (zzr0ck3r):

if its the inverse then h(f(x)) = x = f(h(x))

OpenStudy (zzr0ck3r):

what is h(f(x))?

OpenStudy (anonymous):

f(y) = x = 5/(y - 2) y - 2 = 5/x y = 5/x + 2 And that is the right side of h(x), so it is an inverse.

OpenStudy (anonymous):

Original function:

OpenStudy (anonymous):

inverse:

OpenStudy (anonymous):

i dont see how you got that last conclusion... how did you get y-2 over on the other side without multiplying with x?

OpenStudy (anonymous):

Set the reciprocals equal to each other as a first step. It's faster and not really a short-cut.

OpenStudy (anonymous):

1/x = (y-2)/5

OpenStudy (anonymous):

Then, multiply both sides by 5: 5/x = y - 2 The last step, just add 2 to both sides

OpenStudy (zzr0ck3r):

y=5/(x-2) y(x-2)=5 yx-2y=5 yx=5+2y x=(5+2y)/y thus it is not the inverse...

OpenStudy (anonymous):

@zzr0ck3r you have to switch the "x" and "y" first. Do that, THEN solve for "x" and you'll see that they are inverses.

OpenStudy (zzr0ck3r):

they are not (5+2x)/x is not equal to 5/(x+2)

OpenStudy (anonymous):

They are not supposed to be.

OpenStudy (zzr0ck3r):

I thought you were saying the two function in the question asked are inverses of each other, and they are not.

OpenStudy (anonymous):

I just proved that they are inverses. To find out if functions are inverses, you take the original function, switch the variables and solve. You are forgetting to switch the variables first.

OpenStudy (anonymous):

Look at the 2 functions, defined for x >= 0 y = x^2 and y = +sqrt(x) You will see that they are not equal and are inverses

OpenStudy (zzr0ck3r):

I thought 5/x + 2 = \[\frac{5}{x+2}\].... sorry

OpenStudy (zzr0ck3r):

sorry @tcarroll010

OpenStudy (anonymous):

np! You were eventually able to see it @zzr0ck3r

OpenStudy (zzr0ck3r):

I thought you guys were crazy for a min:)

OpenStudy (anonymous):

What do you mean? I'm crazy a lot! lol

OpenStudy (zzr0ck3r):

I wish I could medal you twice:)

OpenStudy (anonymous):

Another way to see the inverse relationship: 5 / [(5/x + 2) - 2] = x and 5/[5/(x - 2)] + 2 = x

OpenStudy (zzr0ck3r):

right that was my iff statement about f(h(x)) = x = h(f(x))

OpenStudy (anonymous):

yes, I knew you knew about inverses and substituting back. You're way too good not to, so I got a little surprised when you didn't see it right away. I figured you misread the function and that was all.

OpenStudy (zzr0ck3r):

I feel lame:)

OpenStudy (anonymous):

No way! I wish I made as few mistakes as a lot of the others. No lameness! Only good discussions! So, anyway, @nechamoosh are you all set yet?

OpenStudy (anonymous):

oh. yeah sorry! thankss again for the help!!!

OpenStudy (anonymous):

ha, both of you!

OpenStudy (anonymous):

Yes, zzrock3r ' s contribution was definitely worthwhile and had a lot of good theory. His help was very valuable. And I'm glad we all learned something here. I sure did!

OpenStudy (anonymous):

But now, it's Miller time. So, I'll stick around just briefly if you have any last minute questions, @nechamoosh So, fire away if anything is unclear. Otherwise, nice working with you!

OpenStudy (anonymous):

Leaving on a bit of humor, this problem I was "dude". Last problem, I was a lifeguard. And here I thought I was just Sheldon!

OpenStudy (anonymous):

um well there is one last thing i dont get. if A = {l, m, n, o} and B = {10, 20, 30, 40}. is the set of ordered pairs {(l,30) (m,20) (n,10)} a function from set A to set B? why or why not?

OpenStudy (anonymous):

i am completely lost on that one!

OpenStudy (anonymous):

Yes, it's a function because each of the values in "A" (your independent variable set) takes on at most only one value. If, for instance, you had 2 different values to go with "m", like another point being (m, 70), then you wouldn't have a function. It would then be only a relation.

OpenStudy (anonymous):

ohhhhh, okay i get it! thanks so much!!

OpenStudy (anonymous):

Consider the function y = 2x. That's a function because for each "x", there is only one "y". The inverse is also a function, y = x/2 .Example of a function where the inverse is only a relation: y = x^2

OpenStudy (anonymous):

uw! So, if you have a new question, just start it in a new post and someone will come along. If you have any question with what we've covered, then it's ok to ask it here.

OpenStudy (anonymous):

nah im all good! but thanks